The next dividend payment by Hoffman, Inc., will be $2.65 per share. The dividends are anticipated to maintain a growth rate of 4.5 percent forever. If the stock currently sells for $43.15 per share, what is the required return?
Answer #1

Ans 10.64%

P0 = Price of Share
D1 = Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
43.15 = 2.65 / (Ke - 4.5%)
Ke - 4.5% = 2.65 / 43.15
Ke - 4.5% = 6.14%
Ke = 6.14% + 4.5%
Ke = 10.64%